# AlignDiT: Multimodal Aligned Diffusion Transformer for Synchronized Speech Generation

### v1

@@ 5.4 Applications
Video-to-speech. Unlike ADR, video-to-speech takes silent video (without text) as input to read lip movements and synthesize corresponding speech. Some prior works [9, 10] operate in textless manner, while others, such as LipVoicer [74], leverage off-the-shelf lip reading models to guide speech synthesis. For a fair comparison with LipVoicer, we adopt the same lip reading model [46] to obtain pseudo-text labels. Table 6 summarizes the performance comparison between our approach and existing methods. Interestingly, our model significantly outperforms existing methods across all evaluation metrics, including LipVoicer, which is specifically designed for this task and also utilizes an expert lip reading model. This highlights the robustness of our multimodal alignment approach, demonstrating that it generalizes effectively and achieves superior performance even on tasks beyond its primary design.
@@
KWS-Net [53] 262.9ms 42.6%
CTC-based [38] 124.5ms 60.6%
Transpotter [60] 167.3ms 61.8%
DVFA [35] 97.7ms 80.2%
@@
Table 7: Visual forced alignment benchmark.
Visual forced alignment (VFA). Conventionally, the VFA task involves identifying timestamps for each word or phoneme in silent videos. Since our model can synthesize speech given silent video and text input, we bypass direct comparison of silent video and text, and instead leverage these inputs to generate speech signals. We then apply the Montreal Forced Aligner [48] to align the synthesized speech with the corresponding text, thus determining timestamps of each word for VFA task. It is worth noting that we use a single canonical reference speech across all test samples, removing the need for speaker-specific references. Table 7 presents a comparison of alignment performance on the LRS3 dataset. The synthesized speech from our proposed AlignDiT is highly synchronized with the input video, enabling precise forced alignment with text, resulting in significantly better performance compared to existing methods [53, 60, 38, 35] specifically designed for this task. These results support that our model generates highly accurate and temporally synchronized speech from multimodal inputs.
6 Conclusion

### v2

@@ 5.4. Applications
Video-to-speech. Unlike ADR, video-to-speech takes silent video (without text) as input to read lip movements and synthesize corresponding speech. Some prior works (Choi et al., 2023a, b) operate in textless manner, while others, such as LipVoicer (Yemini et al., 2024), leverage off-the-shelf lip reading models to guide speech synthesis. For a fair comparison with LipVoicer, we adopt the same lip reading model (Ma et al., 2023) to obtain pseudo-text labels. Table 6 summarizes the performance comparison between our approach and existing methods. Interestingly, our model significantly outperforms existing methods across all evaluation metrics, including LipVoicer, which is specifically designed for this task and also utilizes an expert lip reading model. This highlights the robustness of our multimodal alignment approach, demonstrating that it generalizes effectively and achieves superior performance even on tasks beyond its primary design.
Table 7. Visual forced alignment benchmark.
@@
KWS-Net (Momeni et al., 2020) 262.9ms 42.6%
CTC-based (Kürzinger et al., 2020) 124.5ms 60.6%
Transpotter (Prajwal et al., 2021) 167.3ms 61.8%
DVFA (Kim et al., 2023) 97.7ms 80.2%
@@
Visual forced alignment (VFA). Conventionally, the VFA task involves identifying timestamps for each word or phoneme in silent videos. Since our model can synthesize speech given silent video and text input, we bypass direct comparison of silent video and text, and instead leverage these inputs to generate speech signals. We then apply the Montreal Forced Aligner (McAuliffe et al., 2017) to align the synthesized speech with the corresponding text, thus determining timestamps of each word for VFA task. It is worth noting that we use a single canonical reference speech across all test samples, removing the need for speaker-specific references. Table 7 presents a comparison of alignment performance on the LRS3 dataset. The synthesized speech from our proposed AlignDiT is highly synchronized with the input video, enabling precise forced alignment with text, resulting in significantly better performance compared to existing methods (Momeni et al., 2020; Prajwal et al., 2021; Kürzinger et al., 2020; Kim et al., 2023) specifically designed for this task. These results support that our model generates highly accurate and temporally synchronized speech from multimodal inputs.
6. Conclusion
@@
Acknowledgements.
This work was supported by IITP grants funded by the Korean government (MSIT, RS-2025-02263169, Detection and Prediction of Emerging and Undiscovered Voice Phishing and RS-2024-00457882, National AI Research Lab Project).